Write an equation and solve for x if the area of the rectangle is 55 square units. the length of a side is 3x-1 and the height is 5
Answer by mananth(16946)

length = 3x-1
height = 5
length * height = 55
(3x-1) *5 = 55
15x-5 =55
15x= 60
x=4 units